(b) Find the impedance of a series RLC circuit if the inductive reactance, capacitive 

Draw the corresponding voltage versus time graph. [10 marks]

reactance and resistance are 184 Ω, 144 Ω and -j30 Ω respectively. Also calculate the 

(a) Write down the equation for a sinusoidal voltage of 50 Hz and its peak value is 20 V. 

phase angle between voltage and current. 


1
Expert's answer
2021-10-07T23:42:26-0400

a)

"V_o=20V\\
f=50Hz\\
V(t)=V_o\sin{2\pi{ft}}\\
V(t)=20\sin100\pi{t}"



b)

"Z=\sqrt{R^2+(X_L-X_C)^2}\\
Z=\sqrt{30^2+(184-144)^2}=50\Omega\\
\phi=\arctan{\dfrac{X_L-X_C}{R}}=\arctan{\dfrac{40}{30}}=53.1\degree"

Impedance is 50 ohms while the phase angle between voltage and current is 53.1degree